# Recurrent Inference Machine for Medical Image Registration

### v1

@@ 2.4 Metrics
$\operatorname{MSE}\left(I_{1},I_{2}\right)=\frac{1}{d_{x}d_{y}d_{z}}{\left\|I_{1}-I_{2}\right\|}_{2}^{2}.$ (14)
The MSE metric is minimized when pixels of $I_{1}$ and $I_{2}$ have the same intensities. Therefore, it is sensitive to the contrast change. In comparison, the NCC metric measures the difference between images with the image intensity normalized. The NCC difference between $I_{1}$ and $I_{2}$ is given by:
$\operatorname{NCC}(I_{1},I_{2})=\frac{1}{d_{x}d_{y}d_{z}}\sum_{\bm{x}\in\Omega_{I_{1}}}\frac{\sum_{\bm{x}^{\prime}\in\Omega_{\bm{x}}}(I_{1}(\bm{x}^{\prime})-\bar{I_{1}}(\bm{x}))(I_{2}(\bm{x}^{\prime})-\bar{I_{2}}(\bm{x}))}{\sqrt{\hat{I}_{1}(\bm{x})\hat{I}_{2}(\bm{x})}},$ (15)
where $\Omega_{I_{1}}$ denotes all possible coordinates in $I_{1}$, $\Omega_{\bm{x}}$ represents a neighborhood of voxels around coordinate position $\bm{x}$ and $\bar{I}(\bm{x})$ and $\hat{I}(\bm{x})$ denote the (local) mean and variance in $\Omega_{\bm{x}}$. In this study, we consider the local NCC with a window size of 9.
@@
Regularization Metrics: To ensure a smooth and reasonable deformation field, we use a diffusion regularization loss which penalizes large displacements in $\bm{\phi}$ acting on $I\in\mathbb{R}^{d_{x}\times d_{y}\times d_{z}}$ (Balakrishnan et al., 2019):
$\mathcal{L}_{\text{reg}}=\frac{1}{d_{x}d_{y}d_{z}}\sum_{\bm{x}\in\Omega_{I}}\|\nabla\bm{\phi}(\bm{x})\|_{2}^{2},$ (17)
where $\nabla\bm{\phi}(\bm{x})$ denotes the Jacobian of $\bm{\phi}$ at coordinate $\bm{x}$. It is noticeable that Eq. 17 and its gradient are not evaluated in each RIIR inference step as indicated in Eq. 10, the outer loss $\mathcal{L}_{\text{outer}}$ and the data-driven training process can guide the RIIR Cell $g_{\theta}$ to learn the regularization implicitly.
@@ 3.1 Dataset
We evaluated our proposed RIIR framework on two separate datasets: 1) A 3D brain MRI image dataset with inter-subject registration setup, OASIS (Marcus et al., 2007) with pre-processing from Hoopes et al. (2021), denoted as OASIS. 2) A 2D quantitative cardiac MRI image datasets based on multiparametric SAturation-recovery single-SHot Acquisition (mSASHA) image time series (Chow et al., 2022), denoted as mSASHA; These datasets, each serving our interests in inter-subject tissue alignment and respiratory motion correction with contrast variation.
@@
where $(TS,TE,TD)$ denotes the set of three acquisition variables, and $(T_{1},T_{2},A)$ is the set of parameters to be estimated for each voxel coordinate of the image series. We encourage interested readers to refer to Chow et al. (2022) for a more detailed explanation.
In our experiment, an in-house mSASHA dataset was used. This fully anonymized raw dataset was provided by NIH, and was considered “non-human subject data research” by the NIH Office of Human Subjects Research”. The dataset comprises 120 subjects, with each subject having 3 slice positions, resulting in a total number of 360 slices. Each mSASHA time series consists of a fixed length of $N=30$ images. We split mSASHA into train/validation/test with counts of $[84,12,24]$ by subjects to avoid data leakage across the three slices. Given variations in image sizes due to different acquisition conditions, we first center-cropped the images into the same size of $144\times 144$. Subsequently, we applied intensity clipping between $(1\%,95\%)$ percentiles to mitigate extreme signal intensities from the chest wall region. We selected the last image in the series as the template $I_{\text{fix}}$, which is a T2-weighted image with the greatest contrast between the myocardium and adjacent blood pool. An illustrative example of mSASHA images can be found in Fig. 3, showing varying contrasts and non-rigid motion across frames.
@@ 3.2 Evaluation Metrics
[…]
For inter-subject brain MRI, two metrics, Dice score and Hausdorff distance (HD) are considered to evaluate the segmentation quality after registration. Given two sets $X\subset M$ and $Y\subset M$, the Dice score, is defined to measure the overlapping of $X$ and $Y$:
@@
where $d(\cdot,\cdot)$ is a metric (2-norm in this work) on $M$ and $d(x,Y):=\inf_{y\in Y}d(x,y)$. As a remark, in this work, we consider the average across all segmentation labels to calculate the Dice score and HD in OASIS instead of only considering major regions.
@@
$\mathcal{D}_{\text{PCA1}}$ $=\sum_{i=1}^{N}\lambda_{i}-\sum_{j=1}^{L}\lambda_{j}=\text{tr}(K)-\sum_{j=1}^{L}\lambda_{j},$ (22)
$\mathcal{D}_{\text{PCA2}}$ $=\sum_{j=1}^{N}j\lambda_{j}.$ (23)
Both metrics were designed to penalize a long-tail distribution of the spectrum of $K$, and $L$ is a hyperparameter regarding the number of parameters of the signal model. For $\mathcal{D}_{\text{PCA1}}$, an ideal scenario would involve all images perfectly aligning with tissue anatomy and the signal model, resulting in a value of 00. Meanwhile, the interpretation of $\mathcal{D}_{\text{PCA2}}$ further emphasizes the tail of the eigenvalues thus enlarging the gaps across experiments.
@@ 3.3 Experimental Settings
We here summarize the main experiments for evaluation and further ablation experiments for RIIR. For all experiments, the main workflow is to register the image series $I$ of length $N$ in a pairwise manner: that is, we first choose a template $I_{\text{fix}}$, and then perform $N$ registrations. When $N=2$, the registration process simplifies to straightforward pairwise registration, as is the situation for OASIS.
@@
We introduce five data-availability scenarios to evaluate the robustness of the models when data availability is limited, on both datasets, which often happens in both research settings and clinical practices as the number of subjects is heavily limited. The training data availability settings in this study were set to $[5\%,10\%,25\%,50\%,100\%]$ for OASIS and mSASHA. It is worth noticing that for limited data availability scenarios, the data used for training remained the same for all models in consideration, and the leave-out test split remained unchanged for all scenarios.
@@
Unlike most related works utilizing the original RIM framework (Lønning et al., 2019; Sabidussi et al., 2021) where two levels of hidden states are considered, we explored the impact of modifying or even turning off hidden states. In our implementation of convolutional GRU, at most two levels of hidden states $\bm{h}_{t}^{1}$ and $\bm{h}_{t}^{2}$ are considered following most of recent works using RIM (Lønning et al., 2019; Sabidussi et al., 2021, 2023). Two hidden states were supposed to have the same channels as in their corresponding convolutional GRU, where in this study, the channels were set to be $[16,16]$ for 3D OASIS and $[32,16]$ for mSASHA after taking the VRAM consumption into account. From this experiment, all experiments were implemented on the validation split for both datasets.
Experiment 3: Choice of Inner Loss Functions $\mathcal{L}_{\text{inner}}$
As detailed in the Methods section, the purpose $\mathcal{L}_{\text{inner}}$ is different from previous iterative registration methods as $\mathcal{L}_{\text{inner}}$ is not proposed as a choice of input of a convolutional recurrent neural network. We evaluated the ablation experiments from now on mSASHA which includes multiple-image registration with motion and varies in contrasts. Since it has been proposed previously that NMI is a more suitable choice of $\mathcal{L}_{\text{sim}}$ for $\mathcal{L}_{\text{outer}}$ in qMRI registration (de Vos et al., 2020) over NCC and MSE, we focused on the choice of $\mathcal{L}_{\text{inner}}$ to investigate the registration performance and the information propagated through the RIIR. That is, in this experiment, we evaluated the choice of $\mathcal{L}_{\text{outer}}$ with MSE, NCC, and NMI with $\mathcal{L}_{\text{outer}}$ as NMI.
Experiment 4: Inclusion of Gradient of Inner Loss $\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}$ as RIIR Input
We performed an experimental study on the input composition for RIIR. As shown in Eq. 12, the goal was to study the data efficiency and the registration performance by incorporating the gradient of $\mathcal{L}_{\text{inner}}$ in RIIR. We could achieve the ablation by changing the input of $g_{\theta}$. A comparison with other input modelling strategies seen in Qiu et al. (2022) against the gradient-based input for $g_{\theta}$ was proposed. Depending on whether the moving image is deformed (explicit) or not (implicit), we ended up with three input compositions:
@@
    RIIR Input: $[\bm{\phi}_{t},\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}]$.
@@
This study aimed to provide insights into the impact of different input compositions on the efficiency of RIIR when data availability varies. We conducted the experiment with low data availability choices ($[5\%,10\%]$) to examine the data efficiency induced by the gradient input.
@@
Experiment 5: RIIR Architecture Ablation Since RIIR is the first attempt to formulate and implement the RIM framework for medical image registration, we performed an extensive ablation study on the network architecture of RIIR, including: 1. number of evaluation steps $T$; 2. choice of loss weights $w_{t}$.
3.4 Implementation Details and Baseline Methods
We implemented the RIIR in the following settings for experiment 1: We used the inference steps $T=6$ and exponential weighting for $w_{t}=10^{\frac{t-1}{T-1}}$. We compared our proposed RIIR with two representative methods in the deep learning paradigm: VoxelMorph (Balakrishnan et al., 2019) and a modification of GraDIRN (Qiu et al., 2022). We conducted an in-depth hyperparameter tuning to compare existing methods and RIIR. For VoxelMorph, we followed a standard setting: a U-Net (Ronneberger et al., 2015) with 5 down-sampling and 5 up-sampling blocks as the backbone. The channels used in each down-sampling block are $\left[16,16,32,32,32\right]$. The implementation of GraDIRN followed the authors’ settings suggested in Qiu et al. (2022) with 2 resolutions and 3 iterations per resolution and the explicit warped images as input. The numbers of parameters are given as follows for RIIR, VoxelMorph, and GraDIRN respectively: 61K, 254K, and 173K. The trade-off parameter $\lambda$ for all models was set to $0.05$ when considering MSE (for OASIS as $\mathcal{L}_{\text{sim}}$) and $0.1$ when considering NMI (for mSASHA) as $\mathcal{L}_{\text{sim}}$. The optimizer of all methods remained the same using Adam (Kingma and Ba, 2015) with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The initial learning rate was set to $8\times 10^{-4}$ for all models. The experiments were performed on an NVIDIA RTX 4090 GPU with a VRAM of 24 GB. We will make our code for RIIR publicly available after acceptance, as well as our implementation of the baseline models.

### v2

@@ 2.4 Metrics
$\operatorname{MSE}\left(I_{1},I_{2}\right)=\frac{1}{d_{x}d_{y}d_{z}}{\left\|I_{1}-I_{2}\right\|}_{2}^{2},$ (14)
where $|\Omega_{I}|=d_{x}d_{y}d_{z}$ denotes the all possible coordinates. The MSE metric is minimized when pixels of $I_{1}$ and $I_{2}$ have the same intensities. Therefore, it is sensitive to the contrast change. In comparison, the NCC metric measures the difference between images with the image intensity normalized. The NCC difference between $I_{1}$ and $I_{2}$ is given by:
$\operatorname{NCC}(I_{1},I_{2})=\frac{1}{|\Omega_{I_{1}}|}\sum_{\bm{\chi}\in\Omega_{I_{1}}}\frac{\sum_{\bm{\chi}^{\prime}\in\Omega_{\bm{\chi}}}(I_{1}(\bm{\chi}^{\prime})-\bar{I_{1}}(\bm{\chi}))(I_{2}(\bm{\chi}^{\prime})-\bar{I_{2}}(\bm{\chi}))}{\sqrt{\hat{I}_{1}(\bm{\chi})\hat{I}_{2}(\bm{\chi})}},$ (15)
where $\Omega_{I_{1}}$ denotes all possible coordinates in $I_{1}$, $\Omega_{\bm{\chi}}$ represents a neighborhood of voxels around coordinate position $\bm{\chi}$ and $\bar{I}(\bm{\chi})$ and $\hat{I}(\bm{\chi})$ denote the (local) mean and variance in $\Omega_{\bm{\chi}}$.
@@
Regularization Metrics: To ensure a smooth and reasonable deformation field, we primarily use a diffusion regularization loss which penalizes large displacements in $\bm{\phi}$ acting on $I\in\mathbb{R}^{d_{x}\times d_{y}\times d_{z}}$ (Fischer and Modersitzki, 2002):
$\mathcal{L}_{\text{diff}}=\frac{1}{|\Omega_{I}|}\sum_{\bm{\chi}\in\Omega_{I}}\left\|\nabla\bm{\phi}(\bm{\chi})\right\|_{2}^{2},$ (17)
where $|\Omega_{I}|=d_{x}d_{y}d_{z}$, $\nabla\bm{\phi}(\bm{\chi})$ denotes the Jacobian of $\bm{\phi}$ at coordinate $\bm{\chi}$. It is noticeable that Eq. 17 and its gradient are not evaluated in each RIIR inference step as indicated in Eq. 8, the outer loss $\mathcal{L}_{\text{outer}}$ and the data-driven training process can guide the RIIR Cell $g_{\theta}$ to learn the regularization implicitly.
Since RIIR aims to learn implicit regularization from data with the outer loss, we also include two additional regularization metrics in our ablation study: curvature regularization and linear elastic regularization with fixed elasticity parameters (Fischer and Modersitzki, 2004). The curvature loss penalizes the second spatial derivatives of the displacement field $\bm{\phi}$, encouraging smoothness in the rate of change of $\bm{\phi}$:
$\mathcal{L}_{\text{curv}}=\frac{1}{|\Omega_{I}|}\sum_{\bm{\chi}\in\Omega_{I}}\sum_{i,j=1}^{3}\left\|\frac{\partial^{2}\bm{\phi}(\bm{\chi})}{\partial\chi_{i}\partial\chi_{j}}\right\|_{2}^{2},$ (18)
where $\frac{\partial^{2}\bm{\phi}(\bm{\chi})}{\partial\chi_{i}\partial\chi_{j}}$ denotes the second-order partial derivative of $\bm{\phi}$ w.r.t. dimensions $\chi_{i}$ and $\chi_{j}$ at coordinates $\bm{\chi}$.
The linear elastic regularization aims to regularize the displacement field by considering both the divergence and strain of the field. Its variational formulation can be described as follows (Fischer and Modersitzki, 2004):
$\mathcal{L}_{\text{elas}}=\frac{1}{|\Omega_{I}|}\sum_{\bm{\chi}\in\Omega_{I}}\left(\frac{\lambda_{e}}{2}\left(\operatorname{div}\bm{\phi}(\bm{\chi})\right)^{2}+\frac{\mu}{4}\sum_{i,j=1}^{3}\left\|\frac{\partial\phi_{i}(\bm{\chi})}{\partial\chi_{j}}+\frac{\partial\phi_{j}(\bm{\chi})}{\partial\chi_{i}}\right\|_{2}^{2}\right),$ (19)
where $\operatorname{div}\bm{\phi}(\bm{\chi})=\sum_{i=1}^{3}\frac{\partial\phi_{i}(\bm{\chi})}{\partial\chi_{i}}$ denotes the divergence of the displacement field, $\lambda_{e}$ and $\mu_{e}$ are Lamé parameters controlling the strength of volumetric and shear deformation respectively. The first term penalizes volume changes, while the second term regularizes shear deformation of the displacement field $\bm{\phi}$.
@@ 3.1 Dataset
We evaluated our proposed RIIR framework on two separate datasets: 1) A 3D brain MRI image dataset with inter-subject registration setup, OASIS (Marcus et al., 2007) with pre-processing from Hoopes et al. (2021), denoted as OASIS. 2) A 3D lung CT dataset with intra-subject registration setup, National Lung Screening Trial (NLST)(Aberle et al., 2011), provided and processed by the Learn2Reg challenge (Hering et al., 2022). This dataset is denoted by NLST. 3) A 2D quantitative cardiac MRI image datasets based on multiparametric SAturation-recovery single-SHot Acquisition (mSASHA) image time series (Chow et al., 2022), denoted as mSASHA. These datasets, each serving our interests in inter-subject tissue alignment and respiratory motion correction with and without contrast variation.
@@
Figure 3: An example of NLST dataset for a single subject with $I_{\text{fix}}$ corresponding to the image captured at inspiratory phase and $I_{\text{mov}}$ corresponding to the image captured at expiratory phase.
@@
NLST: We use the public NLST dataset from the Learn2Reg challenge. The dataset consists of 210 subjects with intra-subject inhale and exhale lung CT images that are affinely pre-aligned. The subjects are split into train/validation/test with counts of $[170,10,30]$. Following the convention, during training, only the inhale and exhale images from the same subject are paired, with the inhale image chosen as $I_{\text{fix}}$ for all pairs. The images were preprocessed and resampled to a size of $224\times 192\times 224$. The keypoints and their correspondences in the lung lobe, along with the lung lobe mask, were provided by the organizers, and obtained by the corrField algorithm (Heinrich et al., 2015) and nnU-Net (Isensee et al., 2021) respectively. An example pair of NLST images is shown in Fig. 3, demonstrating the large deformation required for this task.
@@
where $(TS,TE,TD)$ denotes the set of three acquisition variables, and $(T_{1},T_{2},A)$ is the set of parameters to be estimated for each voxel coordinate of the image series. The sequence for mSASHA consists of a reference image without magnetization preparation, a series of saturation recovery (SR) images, and a series of both SR and $T_{2}$-prepared images. We encourage interested readers to refer to Chow et al. (2022) for a more detailed explanation.
In our experiment, an in-house mSASHA dataset was used. This fully anonymized raw dataset was provided by NIH, and was considered “non-human subject data research” by the NIH Office of Human Subjects Research”. The dataset comprises 120 subjects, with each subject having 3 slice positions, resulting in a total number of 360 slices. Each mSASHA time series consists of a fixed length of $N=30$ images. We split mSASHA into train/validation/test with counts of $[84,12,24]$ by subjects to avoid data leakage across the three slices. Given variations in image sizes due to different acquisition conditions, we first center-cropped the images into the same size of $144\times 144$. Subsequently, we applied intensity clipping between $(1\%,95\%)$ percentiles to mitigate extreme signal intensities from the chest wall region. We selected the last image in the series, i.e., in the $T_{2}$ preparation stage, as the template $I_{\text{fix}}$, which is then a T2-weighted image with the greatest contrast between the myocardium and adjacent blood pool. An illustrative example of mSASHA images can be found in Fig. 4, showing varying contrasts and non-rigid motion across frames.
@@ 3.2 Evaluation Metrics
For evaluating the smoothness of the deformation field, we employ three complementary metrics based on the Jacobian of the deformation $J_{\bm{\phi}}=\nabla\bm{\phi}$: (1) the percentage of negative Jacobian determinant $|J_{\phi}|_{\leq 0}$, which quantifies the proportion of regions exhibiting folding or topology-breaking transformations, (2) the standard deviation of the log-Jacobian determinant $\text{std}(\log|J_{\bm{\phi}}|)$, which characterizes the global variation of volume changes, and (3) the magnitude of the spatial gradient of the Jacobian determinant $\operatorname{mag}(\nabla|J_{\bm{\phi}}|)$, which measures the local rate of change in volume deformation. To allow fair comparisons between methods, we meticulously adjust the regularization parameter $\lambda$ for each baseline to achieve comparable levels of smoothness of transformation. To assess registration accuracy, we utilize structural similarity metrics that are independent of optimization objectives $\mathcal{L}_{\text{sim}}$ and $\mathcal{L}_{\text{reg}}$, with dataset-specific metrics detailed in subsequent sections.
For OASIS, two metrics, Dice score and Hausdorff distance (HD) are considered to evaluate segmentation quality after registration. Given two sets $X\subset M$ and $Y\subset M$, the Dice score, is defined to measure the overlapping of $X$ and $Y$:
@@
where $d(\cdot,\cdot)$ is a metric (2-norm in this work) on $M$ and $d(x,Y):=\inf_{y\in Y}d(x,y)$. As a remark, in this work, we consider the average across all segmentation labels to calculate the Dice score and HD in OASIS instead of only considering the major regions.
For the NLST dataset, we evaluate the registration performance using both lung lobe mask overlap and Target Registration Error (TRE) of the keypoints. For a quantitative evaluation of local registration accuracy, TRE is calculated based on the provided keypoint pairs. Given a set of corresponding keypoint pairs $\{(\bm{p}_{i},\bm{q}_{i})\}_{i=1}^{L}$ where $\bm{p}_{i}$ represents a keypoint in the moving image and $\bm{q}_{i}$ its corresponding point in the fixed image, TRE is defined as:
$\text{TRE}=\frac{1}{L}\sum_{i=1}^{L}\|\bm{\phi}(\bm{p}_{i})-\bm{q}_{i}\|_{2},$ (23)
where $\|\cdot\|_{2}$ denotes the Euclidean distance.
@@
$\mathcal{D}_{\text{PCA1}}$ $=\sum_{i=1}^{N}\sigma_{i}-\sum_{j=1}^{L}\sigma_{j}=\text{tr}(K)-\sum_{j=1}^{L}\sigma_{j},$ (25)
$\mathcal{D}_{\text{PCA2}}$ $=\sum_{j=1}^{N}j\sigma_{j},$ (26)
where $\sigma_{i}$ denotes the $i$-th largest eigenvalue of $K$. Both metrics were designed to penalize a long-tail distribution of the spectrum of $K$, and $L$ is a hyperparameter regarding the number of parameters of the signal model. For $\mathcal{D}_{\text{PCA1}}$, an ideal scenario would involve all images perfectly aligning with tissue anatomy and the signal model, resulting in a value of 00. Meanwhile, the interpretation of $\mathcal{D}_{\text{PCA2}}$ further emphasizes the tail of the eigenvalues, thus enlarging the gaps across experiments.
@@
Table 1: Quantitative comparison of different registration methods with $100\%$ data availability. For each evaluation metric except Par. (number of parameters), Mem. (VRAM consumption), and $T$ (training/inference time), we report mean±$\pm$±std. VRAM consumption indicates the peak GPU memory consumption during training. The training/inference time is profiled for a batch on GPU (excluding data loading time), except for elastix where CPU time is reported. For LapIRN, metrics are reported at level 3. The best performance is shown in bold and marked with * if there is a statistically significant difference ($p<0.05$) from the second best by a two-sided Wilcoxon signed-rank test.
Dataset Method Dice HD(mm) $|J_{\bm{\phi}}|_{\leq 0}(\%)$ std$(\log|J_{\bm{\phi}}|)$ $\operatorname{mag}(\nabla|J_{\bm{\phi}}|)$ Par.(K) Mem.(G) $T$ (s)
OASIS Affine $0.563\pm 0.063$ $8.56\pm 1.19$
elastix $0.696\pm 0.040$ $3.89\pm 0.58$ $0.003\pm 0.002$ $0.254\pm 0.045$ $0.015\pm 0.002$ 17.45
VM $0.729\pm 0.030$ $3.61\pm 0.49$ $0.016\pm 0.031$ $0.295\pm 0.115$ $0.029\pm 0.003$ 320 1.57 0.17/0.14
L2O $0.733\pm 0.032$ $3.62\pm 1.01$ $0.021\pm 0.028$ $0.411\pm 0.132$ $0.035\pm 0.003$ 343 20.52 1.19/0.86
GraDIRN $0.739\pm 0.027$ $3.54\pm 0.45$ $0.012\pm 0.003$ $0.206\pm 0.065$ $0.022\pm 0.002$ 269 16.13 0.57/0.25
LapIRN $0.751\pm 0.026$ $3.50\pm 0.44$ $0.010\pm 0.002$ $0.260\pm 0.051$ $0.034\pm 0.002$ 923 3.78 0.16/0.10
RCVM $0.753\pm 0.025$ $\bm{3.46\pm 0.43}^{*}$ $0.002\pm 0.001$ $0.189\pm 0.043$ $0.027\pm 0.003$ 1920 10.20 0.64/0.54
RIIR $\bm{0.756\pm 0.025}^{*}$ $3.48\pm 0.41$ $0.011\pm 0.009$ $0.264\pm 0.073$ $0.029\pm 0.003$ 436 11.82 0.55/0.27
Dataset Method TRE(mm) Dice $|J_{\bm{\phi}}|_{\leq 0}(\%)$ std$(\log|J_{\bm{\phi}}|)$ $\operatorname{mag}(\nabla|J_{\bm{\phi}}|)$ Par.(K) Mem.(G) $T$(s)
NLST Affine $8.43\pm 3.97$ $0.873\pm 0.041$
elastix $5.01\pm 2.92$ $0.946\pm 0.007$ $000.000\pm 0$ 0.000 $0.160\pm 0.031$ $0.011\pm 0.003$ 14.21
VM $3.99\pm 2.48$ $0.957\pm 0.012$ $0.120\pm 0.162$ $0.677\pm 0.399$ $0.057\pm 0.005$ 320 0.93 0.18/0.15
L2O $2.51\pm 2.96$ $0.961\pm 0.019$ $0.451\pm 0.363$ $1.325\pm 0.554$ $0.068\pm 0.007$ 343 11.81 0.83/0.52
GraDIRN $2.89\pm 2.49$ $0.960\pm 0.020$ $0.114\pm 0.068$ $0.454\pm 0.269$ $0.058\pm 0.005$ 279 9.32 0.37/0.15
LapIRN $2.44\pm 1.60$ $0.967\pm 0.006$ $0.073\pm 0.065$ $0.555\pm 0.240$ $0.058\pm 0.004$ 923 2.18 0.14/0.05
RCVM $2.32\pm 1.35$ $\bm{0.968\pm 0.005}^{*}$ $0.229\pm 0.278$ $0.913\pm 0.511$ $0.063\pm 0.005$ 1920 5.89 0.38/0.27
RIIR $\bm{2.21\pm 1.71}^{*}$ $0.966\pm 0.012$ $0.108\pm 0.259$ $0.568\pm 0.481$ $0.048\pm 0.004$ 436 7.12 0.41/0.18
Dataset Method $\mathcal{D}_{\text{PCA1}}$ $\mathcal{D}_{\text{PCA2}}$ $|J_{\bm{\phi}}|_{\leq 0}(\%)$ std$(\log|J_{\bm{\phi}}|)$ $\operatorname{mag}(\nabla|J_{\bm{\phi}}|)$ Par.(K) Mem.(G) $T$(s)
mSASHA Raw $1.28\pm 0.88$ $46.70\pm 10.40$
elastix $0.40\pm 0.38$ $35.81\pm 4.72$ $0.162\pm 0.005$ $0.551\pm 0.685$ $0.022\pm 0.009$ 2.46
VM $0.46\pm 0.42$ $36.49\pm 5.13$ $0.002\pm 0.001$ $0.210\pm 0.082$ $0.053\pm 0.012$ 79 0.06 0.05/0.01
GraDIRN $0.39\pm 0.38$ $35.85\pm 4.51$ $0.002\pm 0.002$ $0.156\pm 0.043$ $0.040\pm 0.012$ 89 0.21 0.09/0.04
LapIRN $0.34\pm 0.32$ $35.11\pm 4.09$ $0.006\pm 0.002$ $0.248\pm 0.128$ $0.054\pm 0.015$ 309 0.09 0.04/0.01
RCVM $\bm{0.32\pm 0.30}^{*}$ $\bm{34.79\pm 3.83}^{*}$ $0.005\pm 0.002$ $0.275\pm 0.081$ $0.073\pm 0.019$ 589 0.13 0.07/0.02
RIIR $0.36\pm 0.36$ $35.46\pm 4.41$ $0.007\pm 0.002$ $0.312\pm 0.042$ $0.039\pm 0.008$ 148 0.24 0.13/0.05
@@ 3.3 Experimental Settings
We here summarize the main experiments for evaluation and further ablation experiments for RIIR. For all experiments, the main workflow is to register the image series $I$ of length $N$ in a pairwise manner: that is, we first choose a template $I_{\text{fix}}$, and then perform $N$ registrations. When $N=2$, the registration process simplifies to straightforward pairwise registration.
@@
We introduce five data-availability scenarios to evaluate the robustness of the models when data availability is limited, on both datasets, which often happens in both research settings and clinical practices as the number of subjects is heavily limited. The training data availability settings in this study were set to $[5\%,10\%,25\%,50\%,100\%]$ for all datasets. It is worth noticing that for limited data availability scenarios, the data used for training remained the same for all models in consideration, and the leave-out test split remained unchanged for all scenarios.
@@
Unlike most related works utilizing the original RIM framework (Lønning et al., 2019; Sabidussi et al., 2021) where two levels of hidden states are considered, we explored the impact of modifying or even turning off hidden states. In our implementation of convolutional GRU, at most two levels of hidden states $\bm{h}_{t}^{1}$ and $\bm{h}_{t}^{2}$ are considered, following recent works using RIM (Lønning et al., 2019; Sabidussi et al., 2021, 2023). Both hidden states were configured with 32 channels in their corresponding convolutional GRU layers. All experiments in this ablation study were performed in the validation split. We present the results for OASIS and NLST as they represent distinct imaging modalities (brain MRI and lung CT) and registration scenarios.
Experiment 3: Inclusion of Gradient of Inner Loss $\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}$ as RIIR Input
We performed an experimental study on the input composition for RIIR. As shown in Eq. 12, the goal was to study the data efficiency and the registration performance by incorporating the gradient of $\mathcal{L}_{\text{inner}}$ in RIIR. We could achieve ablation by changing the input of $g_{\theta}$. A comparison with other input modeling strategies seen in Qiu et al. (2022) was proposed against the gradient-based input for $g_{\theta}$. Depending on whether the moving image is deformed (explicit) or not (implicit), as well as the original RIM formulation, we ended up with four input compositions:
@@
    RIM Input: $[\bm{\phi}_{t},\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}]$;
  4. 4.
    RIIR Input: $[\bm{\phi}_{t},\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}},I_{\text{mov}}\circ\bm{\phi}_{t},I_{\text{fix}}]$.
@@
This study aimed to provide information on the impact of different input compositions on the efficiency of RIIR when data availability varies. We conducted the experiment with two data availability choices ($[5\%,100\%]$) to examine the data efficiency and other potential influences induced by the gradient input.
Experiment 4: Regularization Analysis
It is known that the regularization metric and weight influence the registration performance. In this experiment, we want to investigate the change in the evaluation metrics. For comparison, we used diffusion, curvature, and elastic regularization functions. For elastic regularization, we set the elastic parameters to be $[\lambda_{e}=540.8,\mu=22.5]$ for OASIS and $[\lambda_{e}=45.33,\mu=8]$ for NLST according to the literature (Kumaresan and Radhakrishnan, 1996; Lai-Fook and Hyatt, 2000; Reithmeir et al., 2024).
@@
Experiment 5: RIIR Architecture Ablation Since RIIR is the first attempt to formulate and implement the RIM framework for medical image registration, we performed an ablation study on the RIIR network architecture for the number of evaluation steps $T$.
3.4 Baseline Methods and Implementation Details
We compared our proposed method to various registration methods that are closely related to our interest. We use the same choice of similarity and regularization functions for all deep learning-based methods in the comparative study for fair comparison, unless otherwise indicated. We used MSE loss for OASIS, NCC with window size $w=5$ for NLST, and NMI with $n=32$ bins for mSASHA. We used diffusion regularization for all three datasets in the comparative study. The methods and applicable hyperparameters are described as follows:
  • Elastix (Klein et al., 2009): An iterative optimization-based registration toolbox. Specifically, we used ITK-Elastix (Ntatsis et al., 2023) in Python. Three resolution levels with third-order B-spline transformation and a grid spacing of four were applied for all datasets.
  • VoxelMorph (Balakrishnan et al., 2019): We used $\lambda=0.02$ for OASIS as in the original paper, $\lambda=0.15$ for NLST, and $\lambda=0.3$ for mSASHA. The channels used in each downsampling encoder block were $\left[16,16,32,32,32\right]$. Two layers of activation with channels $\left[16,16\right]$ after the decoder was used.
  • GraDIRN (Qiu et al., 2022): A multi-resolution multi-step deep learning method that uses explicit similarity loss gradient and dense CNN to produce incremental updates. We followed the original implementation with 3 resolutions and 3 steps per resolution and use the last-step output for loss calculation. The training losses were set to the same as VoxelMorph, with weight parameter $\lambda=0.015$ for OASIS, $\lambda=0.125$ for NLST, and $\lambda=0.25$ for mSASHA.
  • LapIRN (Mok and Chung, 2020): A multi-resolution deep learning method with Laplacian pyramid networks. We followed the original implementation of the displacement version with three-stage training and used multi-level NCC in the paper as similarity loss for OASIS, with $\lambda=0.6$. For NLST and mSASHA, we followed the same loss and weight as for VoxelMorph.
  • Learn-to-Optimise (L2O) (Falta et al., 2022): A multi-step deep learning method that uses additional input modalities, including 3D MIND features (Heinrich et al., 2012) and sampled coordinates for 3D datasets. Constrained by the inherent 3D nature of MIND features, we evaluated L2O on the OASIS and NLST datasets. The original paper uses keypoint supervision only; therefore, we replaced the loss function by the same unsupervised loss as used by other baseline methods while keeping the uniform weight for each step, with $\lambda=0.03$ for OASIS and $\lambda=0.225$ for NLST.
  • Recursive-cascaded VoxelMorph (RCVM) (Zhao et al., 2019):A single-resolution iterative deep learning method that uses cascaded U-Nets and composition of the deformation field to generate the final output. We used the same VoxelMorph backbone and loss functions with $\lambda=0.015$ for OASIS, $\lambda=0.25$ for mSASHA, and $\lambda=0.125$ for NLST.
We implemented the RIIR in the following settings for experiment 1: The backbone network is the same as the VoxelMorph in baseline, with additional ConvGRU in the second level with 32 channels. We used the inference steps $T=6$ and exponential weighting for $w_{t}=10^{\frac{t-1}{T-1}}$. To ensure a similar level of smoothness, the trade-off parameter for RIIR was set to $\lambda=0.0125$ for OASIS, $\lambda=0.125$ for NLST, and $\lambda=0.25$ for mSASHA. The optimizer of all methods remained the same using Adam (Kingma and Ba, 2015) with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The initial learning rate was set to $1\times 10^{-4}$ for all models. For all experiments, the maximum epochs was set to 100 epochs with early stopping if the evaluation metrics does not improve for $10$ epochs. The experiments were performed on an NVIDIA RTX 4090 GPU with a VRAM of 24 GB. The source codes for RIIR, with the implementation of the baseline models and data processing are publicly available111gitlab.tudelft.nl/ai4medicalimaging/riir-public.